QI [10] Let the angle θ be the angle that the vector A makes with the +x-axis, measured counterclockwise from that axis. Find the angle θ for a vector that has the following components. (a) A 3.70 m, A, 1.60 m (b) A, -2.60 m, A, 1.70 m (c) A4.30 m, A-3.00 m (d) 2.90 m , A,--3.10 m
Answer #1

-ι( 1.600 -232850 -ι (. 1.70 = 146.80 -1 ( -3.00 n = 214.go 3.70 m) 4310 -4.30 m -310m-31309* 2.90 m